OpenStudy (anonymous):

So, I have 3 equations that I need to find the greatest common monomial... Could I get some help?

RhondaSommer (rhondasommer):

yep

OpenStudy (anonymous):

\[6m ^{3}x ^{4}+30m ^{5} y ^{2}+45m ^{3} x ^{6} y ^{2}\]

OpenStudy (anonymous):

That's the first one.

RhondaSommer (rhondasommer):

\[3m^3\left(15x^6y^2+10m^2y^2+2x^4\right)\]

RhondaSommer (rhondasommer):

would yo like my steps?

OpenStudy (anonymous):

That would be awesome!

RhondaSommer (rhondasommer):

when i look at you orgigional equation, I understand that they can all be divided out by 3

RhondaSommer (rhondasommer):

so writing it all out

RhondaSommer (rhondasommer):

is simple

RhondaSommer (rhondasommer):

\[6m^3x^4+30m^5y^2+45m^3x^6y^2\]

RhondaSommer (rhondasommer):

3*2*m*m*m*x*x*x*x

RhondaSommer (rhondasommer):

3*2*5*m*m*m*m*m*y*y

RhondaSommer (rhondasommer):

3*3*3*m*m*m*x*x*x*x*x*x*y*y

RhondaSommer (rhondasommer):

in each on the is \[3 m^3\]

RhondaSommer (rhondasommer):

and the rest you put in parenthesis :)

RhondaSommer (rhondasommer):

does that make sense?

OpenStudy (anonymous):

Yes! Thank you so much. Do you have time for two more?

RhondaSommer (rhondasommer):

yeah!

RhondaSommer (rhondasommer):

tyoe it in another equation so i dont have to scroll so much pls

RhondaSommer (rhondasommer):

question*

OpenStudy (anonymous):

\[14x ^{3}y ^{5}z ^{2}-98x ^{2}yz ^{2}+28x ^{2}y ^{2}\]

RhondaSommer (rhondasommer):

\[14x^2y\left(xy^4z^2+2y-7z^2\right)\]

RhondaSommer (rhondasommer):

do you need the steps again?

OpenStudy (anonymous):

No. I got that down. Thanks. One more?

OpenStudy (anonymous):

\[10a ^{3}b ^{6}c ^{5}+20a ^{2}b ^{3}-24b ^{2}c ^{3}\]

RhondaSommer (rhondasommer):

\[\quad 2b^2\left(5a^3b^4c^5+10a^2b-12c^3\right)\]

OpenStudy (anonymous):

Oh my god. I love you so much. You're a lifesaver! Thank you so much!

RhondaSommer (rhondasommer):

your welcome :)